# Finite-Sample Non-Parametric Bounds with an Application to the Causal Effect of Workforce Gender Diversity on Firm Performance

### v1

@@ Appendix C Algorithms
Algorithm 1 Monte-Carlo experiment: single–threshold Manski vs. Hybrid band
1:Fixed inputs
2: Replications B=2, [^0] B=2{,}000; firms $n=50$; panel lengths $T\in\{1,2,5\}$;
3: diversity cut-off $\tau^{\circ}=50\%$ (no grid); overall size $\alpha=0.05$;
4: constant treatment effect $\Delta=4$; two-sided critical value $z_{0.975}=1.96$.
5:
6:Step 0 (once per design)
  1. 1.
    Latent support.
    • Design F: known lower bound $a^{\ast}=0$; draw an oracle sample of size $N_{\text{big}}$ and set $b^{\ast}=\max Y_{it}^{0}+\Delta$.
    • Design G: published support $(a^{\ast},b^{\ast})=(-5,5)$.
    • Designs A–E: oracle sample as above and set $a^{\ast}=\min Y_{it}^{0}$, $b^{\ast}=\max Y_{it}^{0}+\Delta$.
7:
8:for $b=1$ to $B$ do $\triangleright$ replication loop
9:  Draw a fresh baseline $\{Y_{it}^{0},D_{it}\}_{i=1,\dots,n;\,t=1,\dots,T}$.
10:  $Y_{it}=Y_{it}^{0}+\Delta D_{it}$.
11:  Empirical extrema $\widehat{a}=\min Y_{it}^{0}$, $\widehat{b}=\max Y_{it}^{0}$.
[…]
13:  Compute $(L^{\mathrm{M}},U^{\mathrm{M}})$ using $(a,b)$.
14:  if design G then
15:   $(L^{\mathrm{H}},U^{\mathrm{H}})\leftarrow(L^{\mathrm{M}},U^{\mathrm{M}})$
16:  else
17:   $\log C\leftarrow\begin{cases}\log(1/\alpha),&\text{design F (one-sided DKW)},\\ \log(2/\alpha),&\text{designs A--E (two-sided DKW)}.\end{cases}$
18:   $\varepsilon_{n}\leftarrow\sqrt{\log C/(2nT)}$
19:   $(\hat{\sigma}_{L},\hat{\sigma}_{U})\leftarrow$ delta-method SEs
20:   $(L^{\mathrm{H}},U^{\mathrm{H}})\leftarrow\left[L^{\mathrm{M}}-\varepsilon_{n}-z_{0.9875}\hat{\sigma}_{L},\;U^{\mathrm{M}}+\varepsilon_{n}+z_{0.9875}\hat{\sigma}_{U}\right]$
21:  end if
22:  $\text{hit}_{\mathrm{M}}[b]\leftarrow\mathbbm{1}\{\Delta\in[L^{\mathrm{M}},U^{\mathrm{M}}]\}$
23:  $\text{hit}_{\mathrm{H}}[b]\leftarrow\mathbbm{1}\{\Delta\in[L^{\mathrm{H}},U^{\mathrm{H}}]\}$
24:end for
25:Output: empirical coverages
[…]
@@ D.4 Data Clustering by Threshold
(%) Women Overall Pre 01/04/2019 Cyclicals Defensives Growth & Innovation Financials Energy $N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$ $\tau_{5}$ 1298 23930 1004 15214 734 9822 119 4837 199 4645 110 3446 136 984 $\tau_{10}$ 1965 23263 1484 14734 1079 9477 158 4798 317 4527 155 3401 238 882 $\tau_{15}$ 3607 21621 2759 13459 2030 8526 246 4710 652 4192 273 3283 368 752 $\tau_{20}$ 6886 18342 5004 11214 3674 6882 505 4451 1432 3412 578 2978 635 485 $\tau_{25}$ 10902 14326 7685 8533 5538 5018 1102 3854 2272 2572 1068 2488 853 267 $\tau_{30}$ 14969 10259 10219 5999 6976 3580 1965 2991 3130 1714 1784 1772 1012 108 $\tau_{35}$ 18937 6291 12634 3584 8167 2389 2850 2106 3872 972 2830 726 1074 46 $\tau_{40}$ 21853 3375 14332 1886 9027 1529 3708 1248 4441 403 3397 159 1099 21 $\tau_{45}$ 23330 1898 15163 1055 9525 1031 4279 677 4703 141 3526 30 1101 19 $\tau_{50}$ 24102 1126 15549 669 9855 701 4592 364 4799 45 3552 4 1108 12 $\tau_{55}$ 24651 577 15886 332 10113 443 4836 120 4833 11 3553 3 1120 0 $\tau_{60}$ 24952 276 16080 138 10335 221 4910 46 4835 9 3556 0 1120 0 $\tau_{65}$ 25081 147 16135 83 10445 111 4928 28 4836 8 3556 0 1120 0 $\tau_{70}$ 25158 70 16175 43 10491 65 4956 0 4839 5 3556 0 1120 0 $\tau_{75}$ 25177 51 16191 27 10510 46 4956 0 4839 5 3556 0 1120 0 $\tau_{80}$ 25182 46 16195 23 10515 41 4956 0 4839 5 3556 0 1120 0 $\tau_{85}$ 25182 46 16195 23 10515 41 4956 0 4839 5 3556 0 1120 0 $\tau_{90}$ 25182 46 16195 23 10515 41 4956 0 4839 5 3556 0 1120 0 $\tau_{95}$ 25182 46 16195 23 10515 41 4956 0 4839 5 3556 0 1120 0

### v2

@@ Appendix C Algorithms
Algorithm 1 MC: single-threshold Manski vs. Hybrid (compact)
1:Inputs: B=2000B{=}2000, n=50n{=}50, $T\in\{1,2,5\}$, $\tau^{\circ}{=}50\%$, $\alpha{=}0.05$, $\Delta{=}4$; $z_{M}{=}\Phi^{-1}(0.975){=}1.96$, $z_{H}{=}\Phi^{-1}(1-\alpha/4){=}2.24$.
2:Per design (once): obtain $(a^{\ast},b^{\ast})$: G uses (5,5)(-5,5); F uses $(0,\max Y^{0}{+}\Delta)$; A-E use $(\min Y^{0},\max Y^{0}{+}\Delta)$ from a large oracle draw.
3:for $b=1,\dots,B$ do
4:  Generate $(Y^{0}_{it},D_{it})$; set $Y_{it}{=}Y^{0}_{it}{+}\Delta D_{it}$; set $(\hat{a},\hat{b}){=}\big(\min Y^{0},\max Y^{0}\big)$.
5:  Analyst support $(a,b){=}\begin{cases}(a^{\ast},b^{\ast}),&\text{G},\\ (0,\hat{b}),&\text{F},\\ (\hat{a},\hat{b}),&\text{A--E}.\end{cases}$
6:  Split at $\tau^{\circ}$: arms $k\in\{0,1\}$ with sizes $N_{k}$, means $\bar{Y}^{(k)}$, shares $p_{k}$.
7:  Manski: point bounds $L_{M}^{0}{=}p_{1}\bar{Y}^{(1)}+p_{0}a-p_{1}b-p_{0}\bar{Y}^{(0)}$, $U_{M}^{0}{=}p_{1}\bar{Y}^{(1)}+p_{0}b-p_{1}a-p_{0}\bar{Y}^{(0)}$.
8:  Delta SEs $(\hat{\sigma}_{L}^{M},\hat{\sigma}_{U}^{M})$; set $L^{\mathrm{M}}{=}L_{M}^{0}{-}z_{M}\hat{\sigma}_{L}^{M}$, $U^{\mathrm{M}}{=}U_{M}^{0}{+}z_{M}\hat{\sigma}_{U}^{M}$.
9:  Hybrid:
10:  if G then $(L^{\mathrm{H}},U^{\mathrm{H}}){\leftarrow}(L^{\mathrm{M}},U^{\mathrm{M}})$
11:  else
12:   For each arm $k$: set $\epsilon_{k}{=}\begin{cases}\sqrt{\log(8/\alpha)/(2N_{k})},&\text{A,B,E (i.i.d.)},\\ (1{+}4C_{\alpha})\sqrt{2\log(8/\alpha)/N_{k}},&\text{C,D (mixing)},\\ \sqrt{\log(4/\alpha)/(2N_{k})},&\text{F (one-sided upper)}.\end{cases}$ and $r_{k}{=}\min\{\epsilon_{k},\;1/2{-}\epsilon_{k}{-}1/N_{k}\}$.
13:   Endpoints $(L^{(k)},U^{(k)}){=}\begin{cases}\big(\hat{F}_{N_{k}}^{(k),-1}(r_{k}{+}\epsilon_{k}),\ \hat{F}_{N_{k}}^{(k),-1}(1{-}r_{k}{-}\epsilon_{k})\big),&\text{A--E},\\ \big(\lambda,\ \hat{F}_{N_{k}}^{(k),-1}(1{-}r_{k}{-}\epsilon_{k})\big),\ \lambda{=}0,&\text{F}.\end{cases}$
14:   Hybrid point bounds $L_{H}^{0}{=}p_{1}\bar{Y}^{(1)}+p_{0}L^{(1)}-p_{1}U^{(0)}-p_{0}\bar{Y}^{(0)}$, $U_{H}^{0}{=}p_{1}\bar{Y}^{(1)}+p_{0}U^{(1)}-p_{1}L^{(0)}-p_{0}\bar{Y}^{(0)}$.
15:   Delta SEs with gradients $\nabla\mathcal{L}{=}(p_{1},-p_{0},\bar{Y}^{(1)}{-}U^{(0)},\,L^{(1)}{-}\bar{Y}^{(0)})$, $\nabla\mathcal{U}{=}(p_{1},-p_{0},\bar{Y}^{(1)}{-}L^{(0)},\,U^{(1)}{-}\bar{Y}^{(0)})$; set $L^{\mathrm{H}}{=}L_{H}^{0}{-}z_{H}\hat{\sigma}_{L}^{H}$, $U^{\mathrm{H}}{=}U_{H}^{0}{+}z_{H}\hat{\sigma}_{U}^{H}$.
16:  end if
17:  Flags: $\mathrm{hit}_{M}[b]{=}\mathbbm{1}\{\Delta\in[L^{\mathrm{M}},U^{\mathrm{M}}]\}$,  $\mathrm{hit}_{H}[b]{=}\mathbbm{1}\{\Delta\in[L^{\mathrm{H}},U^{\mathrm{H}}]\}$.
18:end for
19:Output: $\widehat{P}_{M}{=}B^{-1}\sum_{b}\mathrm{hit}_{M}[b]$,   $\widehat{P}_{H}{=}B^{-1}\sum_{b}\mathrm{hit}_{H}[b]$.
@@ D.4 Data Clustering by Threshold
(%) Women Overall Pre 01/04/2019 Cyclicals Defensives Growth & Innovation Financials Energy
$N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$ $N_{0}$ $N_{1}$
$\tau_{5}$ 1298 23930 1004 15214 734 9822 119 4837 199 4645 110 3446 136 984
$\tau_{10}$ 1965 23263 1484 14734 1079 9477 158 4798 317 4527 155 3401 238 882
$\tau_{15}$ 3607 21621 2759 13459 2030 8526 246 4710 652 4192 273 3283 368 752
$\tau_{20}$ 6886 18342 5004 11214 3674 6882 505 4451 1432 3412 578 2978 635 485
$\tau_{25}$ 10902 14326 7685 8533 5538 5018 1102 3854 2272 2572 1068 2488 853 267
$\tau_{30}$ 14969 10259 10219 5999 6976 3580 1965 2991 3130 1714 1784 1772 1012 108
$\tau_{35}$ 18937 6291 12634 3584 8167 2389 2850 2106 3872 972 2830 726 1074 46
$\tau_{40}$ 21853 3375 14332 1886 9027 1529 3708 1248 4441 403 3397 159 1099 21
$\tau_{45}$ 23330 1898 15163 1055 9525 1031 4279 677 4703 141 3526 30 1101 19
$\tau_{50}$ 24102 1126 15549 669 9855 701 4592 364 4799 45 3552 4 1108 12
$\tau_{55}$ 24651 577 15886 332 10113 443 4836 120 4833 11 3553 3 1120 0
$\tau_{60}$ 24952 276 16080 138 10335 221 4910 46 4835 9 3556 0 1120 0
$\tau_{65}$ 25081 147 16135 83 10445 111 4928 28 4836 8 3556 0 1120 0
$\tau_{70}$ 25158 70 16175 43 10491 65 4956 0 4839 5 3556 0 1120 0
$\tau_{75}$ 25177 51 16191 27 10510 46 4956 0 4839 5 3556 0 1120 0
$\tau_{80}$ 25182 46 16195 23 10515 41 4956 0 4839 5 3556 0 1120 0
$\tau_{85}$ 25182 46 16195 23 10515 41 4956 0 4839 5 3556 0 1120 0
$\tau_{90}$ 25182 46 16195 23 10515 41 4956 0 4839 5 3556 0 1120 0
$\tau_{95}$ 25182 46 16195 23 10515 41 4956 0 4839 5 3556 0 1120 0